OpenStudy (anonymous):

a bottl3e hold 5 pints of lemonade. how much is this in fluid ounces?

OpenStudy (anonymous):

80

OpenStudy (anonymous):

To do conversions you can set up a conversion equation. In this case the conversion value is 1/16 pints to fluid ounces. You can set up an equation like thi\[\frac{ 5pints }{ 1 }\times \frac{ 16 ounces }{ 1 pint }\] to get the equivalent of 5 pints in ounces. Just multiply the fractions and cancel out the "pints" to get the answer in ounces. in this case it would be 80 ounces. This type of conversion equation can be applied to any conversion by setting up the initial value and then the conversion, with the same type of units on the opposite side of the fraction bar. For example, the way the pints unit is in the equation I gave you.